# Beyond Early-Token Bias: Model-Specific and Language-Specific Position Effects in Multilingual LLMs

### v2

@@ 6.3 Position bias is mostly driven by models
While the “Lost in the Middle” phenomenon holds true overall, all models perform worst in the middle, as shown in Table 5, preferences for the beginning versus the end vary across datasets and models. We observe both model-specific bias (Table 6.3) and language-specific bias (Table 6.3). However, the aggregated results in Table 6 suggest that positional bias is primarily model-driven, as discrepancies across positions are more pronounced between models than across languages. Specifically, DeepSeek-7B, Mistral-7B, and Qwen2.5-7B exhibit a preference for later positions, whereas Llama3.1-8B and Gemma-7B favor earlier ones. The strongest positional effect is observed for DeepSeek-7B and Llama3.1-8B, with a performance difference of approximately 0.1 points between positions.
Table 4: Overall performance across strategies, the worst performance is highlighted in red.
Metric Aligned All Zero No Scores
Accuracy 0.695 0.610 0.721
Table 5: Overall performance across positions, the worst performance is highlighted in red.
Metric TOP MIDDLE BOTTOM
Accuracy 0.684 0.666 0.676
Table 6: Model-wise and Langue-wise position bias, highest accuracies reflect bias and are highlighted in bold-green. Results represent accuracy averaged across all strategies Aligned, All Zero and No Scores presented in Table 7 and Table 8. Position indicates correct context position among all contexts.
Position Dataset Mean accuracy Model Mean accuracy
TOP SQuAD2.0 (en) 0.904 DeepSeek 7B 0.683
MIDDLE 0.894 0.697
BOTTOM 0.891 0.753
TOP MTS-SQuAD (ru) 0.861 Gemma 7B 0.497
MIDDLE 0.861 0.492
BOTTOM 0.873 0.470
TOP MLQA (de) 0.646 Mistral 7B 0.748
MIDDLE 0.615 0.729
BOTTOM 0.624 0.751
TOP MLQA (hi) 0.493 Llama3.1 8B 0.743
MIDDLE 0.452 0.671
BOTTOM 0.469 0.655
TOP MLQA (vi) 0.480 Qwen2.5 7B 0.752
MIDDLE 0.463 0.744
BOTTOM 0.483 0.767
Table 7: Language specific bias with resulting accuracy averaged by models
@@
MIDDLE 0.452 0.379 0.558
BOTTOM 0.486 0.397 0.566
Table 8: Model specific bias with resulting accuracy averaged by datasets
@@
MIDDLE 0.750 0.715 0.767
BOTTOM 0.780 0.745 0.777
@@ 6.4 Positional guidance
Sensitivity to prompt guidance. For all considered languages incorrect relevance scoring leads to significant performance decrease. It varies from $4.3\%$ drop for English (0.918 Aligned vs 0.874 All Zero) to $15.6\%$ for Hindi (0.519 Aligned vs 0.363 All Zero). Similar patterns of positional guidance instability are observed for models. All models except Gemma exhibit pronounced sensitivity to positional cues when contextual scoring is perturbed. The introduction of misleading scores results in a marked decline in accuracy ranging from $1.4\%$ drop for DeepSeek performance (from $0.683$ Aligned to $0.669$ All-Zero), to $25.4\%$ drop for Llama performance (from $0.735$ Aligned to $0.481$ All-Zero). The absence of fluctuations in accuracy for Gemma is probably observed, because of poor quality of this model for all setups and general inability to handle positional guidance. Overall, the majority of setups are sensitive to positional guidance, so such techniques could be utilized to handle position bias.
Score Omission Enhances Robustness. Notably, the No Scores consistently outperforms other strategies. For languages there is a little improvement for Russian in BOTTOM position( $0.889$ Aligned vs $0.888$ No Scores). Similarly, only for Qwen2.5 7B( $0.780$ Aligned vs $0.777$ No Scores) and DeepSeek 7B( $0.761$ Aligned vs $0.756$ No Scores) in BOTTOM position the minor accuracy improvements are observed. For all other setups the performance is deteriorate, ranging from $0.1\%$ to $10.2\%$ accuracy drop for languages and from $0.2\%$ to $7.1\%$ for models (Aligned vs No Scores). This performance degradation is particularly pronounced in low-resource languages such as Vietnamese. These findings challenge previous works (Zhang et al., 2024a) and force to thorough validation of guiding strategies.
@@
Table 9: Predictive Entropy across positions, models and strategies. Highest mean values across strategies are highlighted in red.
Model Position Aligned All Zero No Scores Entropy
@@
MIDDLE 0.106 0.104 0.112 0.107
BOTTOM 0.106 0.101 0.108 0.105
@@ Appendix A RAG systems with context reordering
REAR (Wang et al., 2024a) – integrates document relevance scores into LLMs via embeddings, guiding generation to use internal knowledge (low relevance) or external evidence (high relevance).
Long-Context LLMs Meet RAG (Jin et al., 2024) – addresses the "lost-in-the-middle" problem by reordering retrieved documents, placing highest-scoring ones at sequence boundaries to optimize LLM attention.
OP-RAG (Yu et al., 2024a) – order preserving RAG preserves original document order (vs. sorting chunks), demonstrating improved answer quality through position-aware context organization. However, authors do not mention multidocument scenario.
Clustering & Reordering RAG (Alessio et al., 2024) - cluster sentences by query similarity and sort clusters in descending similarity order for improved retrieval quality.
@@ Appendix B Datasets
SQuAD2.0333https://huggingface.co/datasets/rajpurkar/squad_v2 (Rajpurkar et al., 2018) is an English reading-comprehension benchmark built on Wikipedia passages. It combines 100 000 span-answerable questions from SQuAD 1.1 (Rajpurkar et al., 2016) with more than 50 000 adversarial questions whose answers are deliberately absent.
MTS-SQuAD [^4] https://huggingface.co/datasets/MTS-AI-SearchSkill/MTSBerquad is an extension of the SberQuAD dataset (Efimov et al., 2020) which is a Russian counterpart of SQuAD 2.0. It includes more than 60 000 quesion-answer pairs with improved readability and consistency.
MLQA [^5] https://github.com/facebookresearch/MLQA (Lewis et al., 2020) is a multilingual benchmark built from aligned Wikipedia passages including 12 000 question-answer pairs in English and about 5 000 in each of the other six languages: Arabic, German, Spanish, Hindi, Vietnamese, and Simplified Chinese. Among these languages, we utilized German, Hindi and Vietnamese.
@@ Appendix D Effect of position bias on Entropy
Multilingual Caveat. In multilingual large language models (LLMs), position bias may vary across different layers; for instance, early layers often prioritize English tokens, whereas later layers tend to align more closely with the language of the input prompt (Zhong et al., 2024; Schut et al., 2025). This shift may introduce language-specific positional preferences, thereby challenging the assumption of static position bias. Nevertheless, if homogenization occurs, the overall entropy dynamics remain consistent.
@@ Appendix E Word Order Analysis
In this section, we perform additional investigation if there is the relationship between the position of relevant context, model behavior, and the dominant word order of a language. Specifically, we examine whether position bias amplifies or diminishes the influence of a model’s dominant language word order. Our analysis focuses on the No Scores configuration with five contexts, using Hindi and German as representative languages due to their non-SVO dominant word orders, as documented in The World Atlas of Language Structures. Sentences are parsed using Stanza (Qi et al., 2020), following a methodology similar to that of (Choi et al., 2021). For each verb, we identify its dependents; if a verb has both a subject (indicated by a "subj" relation, or such a relation for the nearest preceding verb connected via a "conj" dependency) and an object (indicated by an "obj" relation), we record the word order of these triplets using the abbreviations "S", "V", and "O". The distribution of "SVO" and "SOV" patterns is then analyzed relative to all extracted triplets.
We find no evidence to suggest that position bias influences models to favor specific word orders. For instance, in Hindi a predominantly subject–object–verb (SOV) language—one might expect subject–verb–object (SVO) rates to be lowest when relevant context appears last for Qwen and first for Llama, however, this pattern is not observed (Table 10). In German, where dominant word order varies by clause type (SVO in main clauses and SOV in subordinate clauses;  (Dryer, 2013)), we examine the prevalence of SVO (quantified as the SVO–SOV difference) in Table 11. Here, also, no association with position bias is detected, as the observed prevalence appears to arise naturally from the proportion of complex sentences.
Model Word order TOP MIDDLE BOTTOM
@@
For model-level analysis, the following distinctions are observed: (1) Llama generates a higher proportion of subject–verb–object (SVO) sentences than Qwen across Hindi, German, and Russian (Table 12)—note that Russian is included despite its typological preference for SVO order due to its rich morphological system, which permits considerable word-order flexibility (Ghosh et al., 2024a); (2) Llama produces syntactically more complex sentences than Qwen across all three contextual positions (Table 13); (3) both models exhibit a tendency to generate more complex sentences when relevant context is provided in the initial position (Table 14).
@@ F.1 LLM-as-a-Judge Verification
Krippendorff’s alpha and Pearson correlation coefficients, calculated for each experimental setup can be seen in Tables 15 and 16 correspondingly. Comparison of human and Llama evaluation can be seen in Figure 3.
@@ Appendix H Non Aggregated results for main experiments
Our non aggregated results for LLM as an accuracy and entropy are presented in Tables 17,18 and 19, 20 correspondingly.
Table 17: Accuracy for three models: DeepSeek 7B Chat, Gemma 7B and Mistral 7B
DeepSeek 7B Chat Gemma 7B Mistral 7B
@@
Table 18: Accuracy for two models: Llama3.1 8B and Qwen2.5 7B
@@
Table 19: Predictive entropy values for three models: DeepSeek 7B Chat, Gemma 7B and Mistral 7B
@@
Table 20: Predictive entropy values for two models Llama3.1 8B and Qwen2.5 7B
@@ Appendix I Prompts
Our user–prompts for LLM-inference in terms of context placement strategies can be seen in Tables 21, 22. For Aligned and All Zero strategies items in a contexts–list has a following format: "- [{score}] {document}". For No Scores strategy items has the following format: "- {document}". As a system–prompt the same instruction was used for all languages (translated correspondingly): "You are an AI assistant who helps solve user issues.".
@@
Language User Prompt
English [Uncaptioned image]
Russian [Uncaptioned image]
German [Uncaptioned image]
Hindi [Uncaptioned image]
Vietnamese [Uncaptioned image]

### v3

@@ 6.3 Position bias is mostly driven by models
While the “Lost in the Middle” phenomenon holds true, all models perform worst in the middle, as shown in Table 4, preferences for the beginning versus the end vary across datasets and models. We observe both model-specific bias (Table 7) and language-specific bias (Table 6). However, the aggregated results in Table 5 suggest that positional bias is primarily model-driven, as discrepancies across positions are more pronounced between models than across languages. Specifically, DeepSeek-7B, Mistral-7B, and Qwen2.5-7B exhibit a preference for later positions, whereas Llama3.1-8B and Gemma-7B favor earlier ones. The strongest positional effect is observed for DeepSeek-7B and Llama3.1-8B, with a performance difference of approximately 0.1 points between positions.
Strategy Position
Aligned All Zero No Scores TOP MID BOT
0.695 0.610 0.721 0.684 0.666 0.676
Table 4: Accuracy score for different Strategies of position-bias elimination(Aligned, All Zero, No Scores) and Positions of relevant context(TOP, MIDDLE, BOTTOM). The results in each column are averaged across all other setups. The worst performance is highlighted in blue.
Position Dataset Acc Model Acc
TOP en 0.904 DeepSeek 7B 0.683
MID 0.894 0.697
BOT 0.891 0.753
TOP ru 0.861 Gemma 7B 0.497
MID 0.861 0.492
BOT 0.873 0.470
TOP de 0.646 Mistral 7B 0.748
MID 0.615 0.729
BOT 0.624 0.751
TOP hi 0.493 Llama3.1 8B 0.743
MID 0.452 0.671
BOT 0.469 0.655
TOP vi 0.480 Qwen2.5-7B 0.752
MID 0.463 0.744
BOT 0.483 0.767
Table 5: Model-wise and Language-wise position bias, highest accuracies reflect bias and are highlighted in bold-green. Results represent accuracy averaged across all strategies Aligned, All Zero and No Scores presented in Table 6 and Table 7. Position indicates correct context position among all contexts.
@@
MID 0.452 0.379 0.558
BOT 0.486 0.397 0.566
Table 6: Language specific bias with resulting accuracy averaged by models.
@@
MID 0.750 0.715 0.767
BOT 0.780 0.745 0.777
Table 7: Model specific bias with resulting accuracy averaged by datasets.
@@ 6.4 Positional guidance
Sensitivity to prompt guidance. For all considered languages incorrect relevance scoring leads to significant performance decrease. It varies from $4.3\%$ drop for English (0.918 Aligned vs 0.874 All Zero) to $15.6\%$ for Hindi (0.519 Aligned vs 0.363 All Zero). All models except Gemma exhibit pronounced sensitivity to positional cues when contextual scoring is perturbed. The introduction of misleading scores results in a marked decline in accuracy ranging from $1.4\%$ drop for DeepSeek performance (from $0.683$ Aligned to $0.669$ All-Zero), to $25.4\%$ drop for Llama performance (from $0.735$ Aligned to $0.481$ All-Zero). The absence of fluctuations in accuracy for Gemma is probably observed, because of poor quality of this model for all setups and general inability to handle positional guidance. Overall, the majority of setups are sensitive to positional guidance, so such techniques could be utilized to handle position bias.
Score Omission Enhances Robustness. Notably, the No Scores consistently outperforms other strategies. For languages there is a little improvement for Russian in BOTTOM position( $0.889$ Aligned vs $0.888$ No Scores). Similarly, only for Qwen2.5 7B( $0.780$ Aligned vs $0.777$ No Scores) and DeepSeek 7B ($0.761$ Aligned vs $0.756$ No Scores) in BOTTOM position the minor accuracy improvements are observed. For all other setups the performance is deteriorate, ranging from $0.1\%$ to $10.2\%$ accuracy drop for languages and from $0.2\%$ to $7.1\%$ for models (Aligned vs No Scores). This performance degradation is particularly pronounced in low-resource languages such as Vietnamese. These findings challenge previous works (zhang2024can) and force to thorough validation of guiding strategies.
@@
Model Position Aligned All Zero No Scores Mean
@@
MID 0.106 0.104 0.112 0.107
BOT 0.106 0.101 0.108 0.105
Table 8: Predictive Entropy - PE across positions, models and strategies. Highest mean values across strategies are highlighted in purple.
@@ Appendix A RAG systems with context reordering
REAR (wang2024rear) – integrates document relevance scores into LLMs via embeddings, guiding generation to use internal knowledge (low relevance) or external evidence (high relevance).
Long-Context LLMs Meet RAG (jin2024long) – addresses the "lost-in-the-middle" problem by reordering retrieved documents, placing highest-scoring ones at sequence boundaries to optimize LLM attention.
OP-RAG (yu2409defense) – order preserving RAG preserves original document order (vs. sorting chunks), demonstrating improved answer quality through position-aware context organization. However, authors do not mention multidocument scenario.
Clustering & Reordering RAG (alessio2024improving) - cluster sentences by query similarity and sort clusters in descending similarity order for improved retrieval quality.
@@ Appendix B Datasets
SQuAD2.0333https://huggingface.co/datasets/rajpurkar/squad_v2 (rajpurkar2018know) is an English reading-comprehension benchmark built on Wikipedia passages. It combines 100 000 span-answerable questions from SQuAD 1.1 (rajpurkar2016squad) with more than 50 000 adversarial questions whose answers are deliberately absent.
MTS-SQuAD [^4] https://huggingface.co/datasets/MTS-AI-SearchSkill/MTSBerquad is an extension of the SberQuAD dataset (efimov2020sberquad) which is a Russian counterpart of SQuAD 2.0. It includes more than 60 000 quesion-answer pairs with improved readability and consistency.
MLQA [^5] https://github.com/facebookresearch/MLQA (lewis2020mlqa) is a multilingual benchmark built from aligned Wikipedia passages including 12 000 question-answer pairs in English and about 5 000 in each of the other six languages: Arabic, German, Spanish, Hindi, Vietnamese, and Simplified Chinese. Among these languages, we utilized German, Hindi and Vietnamese.
@@ Appendix D Effect of position bias on Entropy
Multilingual Caveat. In multilingual large language models (LLMs), position bias may vary across different layers; for instance, early layers often prioritize English tokens, whereas later layers tend to align more closely with the language of the input prompt (zhong2024beyond; schut2025multilingual). This shift may introduce language-specific positional preferences, thereby challenging the assumption of static position bias. Nevertheless, if homogenization occurs, the overall entropy dynamics remain consistent.
@@ Appendix E Word Order Analysis
In this section, we perform additional investigation if there is the relationship between the position of relevant context, model behavior, and the dominant word order of a language. Specifically, we examine whether position bias amplifies or diminishes the influence of a model’s dominant language word order. Our analysis focuses on the No Scores configuration with five contexts, using Hindi and German as representative languages due to their non-SVO dominant word orders, as documented in The World Atlas of Language Structures. Sentences are parsed using Stanza (qi2020stanza), following a methodology similar to that of (choi-etal-2021-investigating). For each verb, we identify its dependents; if a verb has both a subject (indicated by a "subj" relation, or such a relation for the nearest preceding verb connected via a "conj" dependency) and an object (indicated by an "obj" relation), we record the word order of these triplets using the abbreviations "S", "V", and "O". The distribution of "SVO" and "SOV" patterns is then analyzed relative to all extracted triplets.
We find no evidence to suggest that position bias influences models to favor specific word orders. For instance, in Hindi a predominantly subject–object–verb (SOV) language—one might expect subject–verb–object (SVO) rates to be lowest when relevant context appears last for Qwen and first for Llama, however, this pattern is not observed (Table 9). In German, where dominant word order varies by clause type (SVO in main clauses and SOV in subordinate clauses;  (wals-81)), we examine the prevalence of SVO (quantified as the SVO–SOV difference) in Table 10. Here, also, no association with position bias is detected, as the observed prevalence appears to arise naturally from the proportion of complex sentences.
Model Word order TOP MID BOT
@@
For model-level analysis, the following distinctions are observed: (1) Llama generates a higher proportion of subject–verb–object (SVO) sentences than Qwen across Hindi, German, and Russian (Table 11)—note that Russian is included despite its typological preference for SVO order due to its rich morphological system, which permits considerable word-order flexibility (ghosh-etal-2024-morphology); (2) Llama produces syntactically more complex sentences than Qwen across all three contextual positions (Table 12); (3) both models exhibit a tendency to generate more complex sentences when relevant context is provided in the initial position (Table 13).
@@ F.1 LLM-as-a-Judge Verification
Krippendorff’s alpha and Pearson correlation coefficients, calculated for each experimental setup can be seen in Tables 14 and 15 correspondingly. Comparison of human and Llama evaluation can be seen in Figure 3.
@@ Appendix H Non Aggregated results for main experiments
Our non aggregated results for LLM as an accuracy and entropy are presented in Tables 16,17 and 18, 19 correspondingly.
DeepSeek 7B Gemma 7B Mistral 7B
@@
Table 16: Accuracy for three models: DeepSeek 7B, Gemma 7B and Mistral 7B, DeepSeek consistently failed in Hindi
@@
Table 17: Accuracy for two models: Llama3.1 8B and Qwen2.5 7B
@@
Table 18: Predictive entropy values for three models: DeepSeek 7B Chat, Gemma 7B and Mistral 7B, Deep Seek consistently failed in Hindi
@@
Table 19: Predictive entropy values for two models Llama3.1 8B and Qwen2.5 7B.
@@ Appendix I Prompts
Our user–prompts for LLM-inference in terms of context placement strategies can be seen in Tables 20, 21. For Aligned and All Zero strategies items in a contexts–list has a following format: "- [{score}] {document}". For No Scores strategy items has the following format: "- {document}". As a system–prompt the same instruction was used for all languages (translated correspondingly): "You are an AI assistant who helps solve user issues.".
Language User Prompt
English [Uncaptioned image]
Russian [Uncaptioned image]
German [Uncaptioned image]
Hindi [Uncaptioned image]
Vietnamese [Uncaptioned image]